# Self-Supervised Frameworks for Speaker Verification via Bootstrapped Positive Sampling

### v2

@@ III-D SSPS-NN: Nearest Neighbors sampling
Following the k-nearest neighbors (k-nn) algorithm, samples that are close to the utterance in the latent space can be considered as appropriate pseudo-positives.
@@
$\mathcal{N}_{i}\triangleq\operatorname*{top\,k}_{j\neq i}\big{(}\left\{\operatorname{sim}\left(\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),\forall{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}j\in\mathcal{I}}\right\};M\big{)},$ (6)
@@
This approach is similar to NNCLR [dwibedi2021NNCLR] and GPS-SSL [feizi2024GPSSSL] but is limited in the context of speaker recognition because pseudo-positives will most likely belong to the same recording as their anchor, even if the sampling window $M$ is large.
@@ III-E SSPS-Clustering: Clustering-based sampling
Clustering methods can also be employed to sample pseudo-positives from the same class or a neighboring class, given the clustering assignment of the utterance.
@@
The k-means algorithm is applied to the reference representations from $\hat{\bm{Q}}$ to group the ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\hat{\bm{Q}}|}$ utterances into $K$ clusters. The clustering is performed at the beginning of each training epoch to refine assignments as the quality of SSL representations improves. Let $c_{i}$ be the assigned cluster of the $i$-th utterance, and $\bm{m}_{k}$ the centroid of the $k$-th cluster.
@@ III-E1 Same-cluster sampling
Samples belonging to the same cluster as the utterance can be used to select the corresponding pseudo-positives because the clustering groups speaker identities when using an appropriate number of clusters, i.e., the hyperparameter $K$ should tend to the actual number of speakers in the training set. Therefore, the cluster $\hat{c}_{i}$ from which the pseudo-positive of the $i$-th sample will be selected can be set to the same as the anchor, such that:
@@ III-E2 Neighboring-clusters sampling
Based on the assumption regarding the modeling of source information before speaker identities, samples from neighboring clusters can be considered as pseudo-positives when choosing a number of clusters $K$ that is significantly larger compared to the number of speakers and tends to the number of recordings in the training set. The underlying purpose is to select a class close enough to $c_{i}$ in the representation space to maximize the probability of being from the same speaker while having different channel characteristics. Hence, the sampling cluster $\hat{c}_{i}$ can be defined as:
@@
where $\mathcal{C}_{k}$ is the set of $M$ nearest clusters from the $k$-th cluster in the representation space and is defined as:
@@
$\mathcal{S}_{k}\triangleq\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}j\in\mathcal{I}}\text{ s.t. }c_{j}=k\}.$ (12)
@@
The preliminary experiments also include SSPS-Clustering (C), which corresponds to using $\bm{m}_{\hat{c}_{i}}$ (i.e., the centroid of the sampling cluster $\hat{c}_{i}$) instead of $\bm{q}^{\prime}_{pos(i)}$ as the positive embedding.
@@
1:Inputs: ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{B}}$, $\hat{\bm{Y}}$, $\bm{Z}$, $\bm{Z}^{\prime}$, $\mathcal{L}$, $\bm{Q}^{\prime}$
2:$\bm{Z}_{\text{SSPS}}^{\prime}\leftarrow\left\{\begin{array}[]{ll}\bm{q}^{\prime}_{pos(i)},&\text{if }pos(i)\neq\varnothing,\\ \bm{z}_{i}^{\prime},&\text{otherwise}\end{array}\right\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$triangleright by Eq. (11)
@@ IV-A Datasets and feature extraction
All models are trained on the VoxCeleb2 [chung2018VoxCeleb2] dev set, which was collected ”in the wild” from YouTube videos and contains 1,092,009 utterances from 5,994 speakers distributed within 145,569 videos. The evaluation is performed using VoxCeleb [nagrani2017VoxCeleb] original (VoxCeleb1-O), extended (VoxCeleb1-E) and hard (VoxCeleb1-H) trials. Speaker labels are discarded for the self-supervised training.
The duration of audio frames is 2 seconds by default. For DINO, local and global frames correspond to four 2-second and two 4-second segments, respectively. The length of the reference frame used by the SSPS framework is 4 seconds. Audio segments are randomly sampled from an utterance and may overlap. Input features are obtained by computing 40-dimensional log-mel spectrogram features using the torchaudio library. The Hamming window length is 25 ms, and the frame-shift is set to 10 ms. Voice Activity Detection (VAD) is not applied as training samples consist mostly of continuous speech segments. Input features are normalized using instance normalization.
@@ IV-B Data-augmentation
SSL commonly relies on extensive data-augmentation techniques to learn representations that are robust against extrinsic variabilities such as environmental and channel noise, mismatching recording devices, and varying acoustic conditions.
Following other works on SSL for SV [zhang2021SimCLR, xia2021MoCo, lepage2022LabelEfficientSSLSV], two types of transformations are applied to the input signals at the beginning of each training iteration. First, reverberation is applied by randomly selecting an RIR from the simulated Room Impulse Response database [ko2017StudyReverberantSpeechRobustSR]. Then, the MUSAN dataset [snyder2015MUSAN] is used to randomly sample background noises, overlapping music tracks, or speech segments. To simulate a variety of real-world scenarios, the Signal-to-Noise Ratio (SNR) is randomly sampled between $\left[13;20\right]$ dB for speech, $\left[5;15\right]$ dB for music, and $\left[0;15\right]$ dB for noises.
@@ IV-C SSL frameworks
The encoder $f$ is based on the ECAPA-TDNN [desplanques2020ECAPATDNN] (22.5M parameters) architecture. For preliminary experiments, the encoder $f$ is replaced by the Fast ResNet-34 [chung2020DefenceMetricLearningSR] (1.4M parameters) architecture, enabling faster experimental iteration for SSPS hyperparameters tuning. The Fast ResNet-34 has a base dimension of 16 and relies on Self-Attentive Pooling (SAP) to produce utterance-level representations. The ECAPA-TDNN has a hidden dimension set to 1024 and uses Attentive Statistics Pooling (ASP). Both encoders output representations of dimension $D_{\text{repr}}=512$, which are used to perform SV.
The projector $g$ is a standard MLP composed of three linear layers. A Batch Normalization and an ReLU activation function follow each layer except the last one. The hidden dimension is set to 2048, and the last layer outputs $D_{\text{emb}}=512$ units. The projector is discarded for contrastive methods (SimCLR and MoCo) as it degrades the downstream performance.
Models are trained during 100 epochs using Adam optimizer with no weight decay, a batch size of 256, and a learning rate set to 0.001 and reduced by 5% every 5 epochs. For DINO, the number of epochs is reduced to 80, the optimizer is SGD, the batch size is set to 128, the weight decay is set to $5e^{-5}$, and the learning rate is linearly ramped up to 0.2 during the 10-epochs warm-up before following a cosine scheduler from 0.2 to $1e^{-5}$.
The supervised baseline corresponds to an equivalent model trained with the AAM-Softmax loss ($s=30$, $m=0.2$) in a fully supervised way using the train set labels.
@@ IV-C1 SimCLR
The loss employed is the symmetric version of NT-Xent, and the temperature is set to $\tau=0.03$.
@@ IV-C2 MoCo
The temperature is set to $\tau=0.03$. The memory queue has a size of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\bm{Q}^{\text{MoCo}}|}=65{,}536$ negatives. The coefficient of the EMA update is fixed to $m=0.999$.
@@ IV-C3 SwAV
The temperature is set to $\tau=0.1$ and the number of prototypes to $P=3000$. The queue is enabled after 15 epochs, and its size is set to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\bm{Q}^{\text{SwAV}}|}=3{,}840$. The prototypes are frozen during the first epoch.
@@ IV-C5 DINO
The projector output dimension is reduced to 256, and a final layer is introduced to map the $l_{2}$-normalized embeddings to $D_{\text{emb}}=65{,}536$ units. This last layer has weight normalization and is frozen during the first epoch. The student and teacher temperatures are set to $\tau_{\text{s}}=0.1$ and $\tau_{\text{t}}=0.04$, respectively. The coefficient of the EMA update goes from $m=0.996$ to $m=1.0$ using a cosine scheduler. Gradients with a norm greater than 3.0 are clipped.
@@ IV-D SSPS
Different positive sampling strategies are presented in the experiments: ‘SSL’ corresponds to the default same-utterance sampling used in SSL frameworks, ‘SSPS’ refers to the proposed approach, and ‘Supervised’ represents a supervised positive sampling baseline that generates anchor-positive pairs from different recordings using the train set labels. These systems are trained using the corresponding positive sampling technique during 20 epochs starting from the end of the standard SSL training.
The SSPS algorithm is vectorized and processes mini-batches at each training iteration in Distributed Data-Parallel (DDP) mode. The size of the reference memory queue $\hat{\bm{Q}}$ is set to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\hat{\bm{Q}}|}=N$. The size of the positive memory queue $\bm{Q}^{\prime}$ is set to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\bm{Q}^{\prime}|}=N$ for SSPS-NN and to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\bm{Q}^{\prime}|}=K$ for SSPS-Clustering. K-means is performed at the beginning of every epoch using a custom PyTorch implementation running on GPU for 10 iterations. Experiments are conducted on 2 $\times$ NVIDIA Tesla A100 80 GB and 4 $\times$ NVIDIA Tesla A100 80 GB for DINO.
The hyperparameter search is conducted using the SimCLR framework with the Fast ResNet-34 encoder, and the number of additional training epochs is reduced from 20 to 10. By default, the number of prototypes is set to $K=25{,}000$, and the sampling window hyperparameter is set to $M=50$ and $M=1$ for SSPS-NN and SSPS-Clustering, respectively.
Additional metrics are reported when SSPS is enabled by using the train set labels:
  • Speaker accuracy (%): corresponds to the average number of pseudo-positives samples matching the anchor speaker identity;
  • Recording accuracy (%): corresponds to the average number of pseudo-positives matching the anchor recording (i.e., VoxCeleb video).
The objective of the proposed approach is to simultaneously maximize speaker accuracy while minimizing recording accuracy to encourage diversity in recording conditions between the anchor and pseudo-positive.
@@ IV-E Evaluation protocol
To evaluate the performance of the systems on speaker verification, the model weights of the last 10 epochs are averaged to provide more consistent results and extract representations by processing full-length audio samples from the test set. Then, to determine the scoring of each trial, the cosine similarity of $l_{2}$-normalized embedding pairs is computed.
Following NIST Speaker Recognition evaluation protocol [sadjadi2017NISTSRE2016], the performance is reported in terms of Equal Error Rate (EER) and minimum Detection Cost Function (minDCF) with $P_{target}=0.01$, $C_{\text{miss}}=1$ and $C_{\text{fa}}=1$.
@@ V-A Effect of the positive sampling on SSL performance
Method Positive sampling VoxCeleb1-O $\Delta$ (%)
@@
The supervised baseline achieves an EER of 1.34% on VoxCeleb1-O, highlighting the performance gap that still exists between SSL methods and supervised systems. This work hypothesizes that one of the primary bottlenecks is the same-utterance positive sampling used in SSL frameworks. As demonstrated in Table I, significant performance improvements are observed when leveraging VoxCeleb metadata to create anchor-positive pairs from distinct recordings of the same speaker. This supervised positive sampling results in EER reductions of approximately 73%, 72%, 45%, 41%, and 23% for SimCLR, MoCo, SwAV, VICReg, and DINO, respectively, on VoxCeleb1-O.
These findings reveal that the default SSL same-utterance positive sampling strategy is a major limitation that hinders the performance of SSL frameworks. Adopting a different positive sampling method could significantly enhance the effectiveness of SSL frameworks on SV tasks.
@@ V-B Hyperparameter tuning for SSPS
TABLE II: Hyper-parameters search for SSPS-NN and SSPS-Clustering in terms of SV performance. The model is SimCLR, the encoder is Fast ResNet-34, and the benchmark is VoxCeleb1-O. Best results for each method are in bold.
Positive sampling SSPS Hyper-params. VoxCeleb1-O
@@
Figure 4: Hyperparameters search for SSPS-NN (a) and SSPS-Clustering (b) in terms of EER, pseudo-positives speaker accuracy and pseudo-positives recording accuracy. The model is SimCLR, the encoder is Fast ResNet-34 and the EER is reported on VoxCeleb1-O.
In this work, it is hypothesized that the main bottleneck of SSL frameworks revolves around the way positives are sampled. The preliminary results of SSPS on SV (VoxCeleb1-O), obtained with SimCLR and the Fast ResNet-34 encoder, are reported in Table II with different sampling algorithms and hyperparameters. These results are compared with SSL and Supervised positive sampling techniques, achieving 9.05% EER and 3.93% EER on VoxCeleb1-O. Additionally, Figure 4 illustrates the effect of $K$ and $M$ in terms of EER on VoxCeleb1-O, speaker accuracy, and recording accuracy.

### v3

@@ III-D SSPS-NN: Nearest Neighbors sampling
Following the k-nearest neighbors (k-nn) algorithm, samples that are close to an utterance in the latent space can be considered as appropriate pseudo-positives.
@@
$\mathcal{N}_{i}\triangleq\operatorname*{top\,k}_{j\neq i}\big{(}\left\{\operatorname{sim}\left(\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),\forall j\in\mathcal{I}\right\};M\big{)},$ (6)
@@
This approach is similar to NNCLR [25] and GPS-SSL [26] but is limited in the context of SR because pseudo-positives will most likely belong to the same recording as their anchor, even if the sampling window $M$ is large.
@@ III-E SSPS-Clustering: Clustering-based sampling
Clustering methods can also be employed to sample pseudo-positives from the same class or a neighboring class, given the clustering assignment of an utterance.
@@
The k-means algorithm is applied to the reference representations from $\hat{\bm{Q}}$ to group the $|\hat{\bm{Q}}|$ utterances into $K$ clusters. The clustering is performed at the beginning of each training epoch to refine assignments as the quality of SSL representations improves. Let $c_{i}$ be the assigned cluster of the $i$-th utterance, and $\bm{m}_{k}$ the centroid of the $k$-th cluster.
@@ III-E1 Same-cluster sampling
Samples belonging to the same cluster as an utterance can be used to select the corresponding pseudo-positives because the clustering groups speaker identities when using an appropriate number of clusters, i.e., the hyperparameter $K$ should tend to the actual number of speakers in the training set. Therefore, the cluster $\hat{c}_{i}$ from which the pseudo-positive of the $i$-th sample will be selected can be set to the assigned cluster of this utterance, such that:
@@ III-E2 Neighboring-clusters sampling
Based on the assumption regarding the modeling of recording information before speaker identities, samples from neighboring clusters can be considered as pseudo-positives when choosing a number of clusters $K$ that is significantly larger than the number of speakers and tends to the number of recordings in the training set. The purpose is to select a class close enough to $c_{i}$ in the latent space to maximize the probability of being from the same speaker while having different channel characteristics. Hence, the sampling cluster $\hat{c}_{i}$ can be defined as:
@@
where $\mathcal{C}_{k}$ is the set of $M$ nearest clusters from the $k$-th cluster in the latent space and is defined as:
@@
$\mathcal{S}_{k}\triangleq\{j\in\mathcal{I}\text{ s.t. }c_{j}=k\}.$ (12)
@@
The preliminary experiments also include SSPS-Clustering (C), which corresponds to using $\bm{m}_{\hat{c}_{i}}$ (i.e., the centroid of the sampling cluster $\hat{c}_{i}$) instead of $\bm{q}^{\prime}_{pos(i)}$ as the pseudo-positive embedding.
@@
1:Inputs: $\mathcal{B}$, $\hat{\bm{Y}}$, $\bm{Z}$, $\bm{Z}^{\prime}$, $\mathcal{L}$, $\bm{Q}^{\prime}$
2:$\bm{Z}_{\text{SSPS}}^{\prime}\leftarrow\left\{\begin{array}[]{ll}\bm{q}^{\prime}_{pos(i)},&\text{if }pos(i)\neq\varnothing,\\ \bm{z}_{i}^{\prime},&\text{otherwise}\end{array}\right\}_{i\in\mathcal{B}}$triangleright by Eq. (11)
@@ IV-A Datasets and feature extraction
All models are trained on the VoxCeleb2 [5] dev set, which was collected ”in the wild” from YouTube videos and contains 1 092 00910920091\,092\,0091 092 009 utterances from $5994$ speakers distributed within 145 569145569145\,569145 569 videos. The evaluation is performed using VoxCeleb [34] original (VoxCeleb1-O), extended (VoxCeleb1-E) and hard (VoxCeleb1-H) trials. Speaker labels are discarded for the self-supervised training.
[…]
@@ IV-B Data-augmentation
SSL commonly relies on extensive data-augmentation techniques to learn representations that are robust against extrinsic variabilities, such as environmental and channel noise, recording devices, and varying acoustic conditions.
Following other SSL approaches for SV [15, 16, 35], two types of transformations are applied to the input signals at the beginning of each training iteration. First, reverberation is applied by randomly selecting an RIR from the simulated Room Impulse Response database [36]. Then, the MUSAN dataset [37] is used to randomly sample background noises, overlapping music tracks, or speech segments. To simulate a variety of real-world scenarios, the Signal-to-Noise Ratio (SNR) is randomly sampled between $\left[13;20\right]$ dB for speech, $\left[5;15\right]$ dB for music, and $\left[0;15\right]$ dB for noises.
@@ IV-C SSL frameworks
The encoder $f$ is based on the ECAPA-TDNN [4] (22.5M parameters) architecture. For preliminary experiments, the encoder $f$ is replaced by the Fast ResNet-34 [3] (1.4M parameters) architecture, enabling faster experimental iteration for SSPS hyperparameter tuning. The Fast ResNet-34 has a base dimension of 16 and relies on Self-Attentive Pooling (SAP) to produce utterance-level representations. The ECAPA-TDNN has a hidden dimension set to 1024 and uses Attentive Statistics Pooling (ASP). Both encoders output representations of dimension $D_{\text{repr}}=512$, which are used to perform SV.
The projector $g$ is a standard MLP composed of three linear layers. A Batch Normalization and a ReLU activation function follow each layer except the last one. The hidden dimension is set to 2048, and the last layer outputs $D_{\text{emb}}=512$ units. The projector is discarded for contrastive methods (SimCLR and MoCo) as it degrades the downstream performance.
Models are trained for 100 epochs using Adam optimizer with no weight decay, a batch size of 256, and a learning rate set to 0.001, which is reduced by 5% every 5 epochs. For DINO, the number of epochs is reduced to 80, the optimizer is SGD, the batch size is set to 128, the weight decay is set to $5e^{-5}$, and the learning rate is linearly ramped up to 0.2 during the 10-epochs warm-up before following a cosine scheduler from 0.2 to $1e^{-5}$.
The supervised baseline corresponds to an equivalent model trained with the AAM-Softmax loss ($s=30$, $m=0.2$) in a fully supervised way using the training set labels.
@@ IV-C1 SimCLR
The temperature is set to $\tau=0.03$, and the loss implementation is symmetric.
@@ IV-C2 MoCo
The temperature is set to $\tau=0.03$. The memory queue has a size of $|\bm{Q}^{\text{MoCo}}|=$65\,536$$ negatives. The coefficient of the EMA update is fixed to $m=0.999$.
@@ IV-C3 SwAV
The temperature is set to $\tau=0.1$ and the number of prototypes to $P=3000$. The queue is enabled after 15 epochs, and its size is set to $|\bm{Q}^{\text{SwAV}}|=$3840$$. The prototypes are frozen during the first epoch. The loss implementation is symmetric.
@@ IV-C5 DINO
The projector output dimension is reduced to 256, and a final layer is introduced to map the $l_{2}$-normalized embeddings to $D_{\text{emb}}=$65\,536$$ units. This last layer has weight normalization and is frozen during the first epoch. The student and teacher temperatures are set to $\tau_{\text{s}}=0.1$ and $\tau_{\text{t}}=0.04$, respectively. The coefficient of the EMA update increases from $m=0.996$ to $m=1.0$ using a cosine scheduler. Gradients with a norm greater than 3.0 are clipped.
@@ IV-D SSPS
Different positive sampling strategies are presented in the experiments: ‘SSL’ corresponds to the default same-utterance sampling used in SSL frameworks, ‘SSPS’ refers to the proposed approach, and ‘Supervised’ represents a supervised positive sampling baseline that generates anchor-positive pairs from different recordings using the training set labels. These systems are trained using the corresponding positive sampling technique during 20 epochs starting from the end of the standard SSL training.
The SSPS algorithm is vectorized and processes mini-batches at each training iteration in Distributed Data-Parallel (DDP) mode. The size of the reference memory queue $\hat{\bm{Q}}$ is set to $|\hat{\bm{Q}}|=N$. The size of the positive memory queue $\bm{Q}^{\prime}$ is set to $|\bm{Q}^{\prime}|=N$ for SSPS-NN and to $|\bm{Q}^{\prime}|=K$ for SSPS-Clustering. K-means is performed at the beginning of every epoch using a custom PyTorch implementation running on the GPU for 10 iterations. Experiments are conducted on 2 $\times$ NVIDIA Tesla A100 80 GB and 4 $\times$ NVIDIA Tesla A100 80 GB for DINO.
[…]
Additional metrics are reported when SSPS is enabled by using the training set labels:
  • Speaker accuracy (%): corresponds to the average number of pseudo-positive samples matching the anchor’s speaker identity;
  • Recording accuracy (%): corresponds to the average number of pseudo-positive samples matching the anchor’s recording (i.e., VoxCeleb video).
The objective of the proposed approach is to simultaneously maximize speaker accuracy while minimizing recording accuracy to encourage diversity in recording conditions across anchors and their corresponding pseudo-positives.
@@ IV-E Evaluation protocol
To evaluate the performance of the systems on SV, the model weights of the last 10 epochs are averaged to provide more consistent results. Representations are extracted by processing full-length audio samples from the test set. The scoring of each trial is determined by computing the cosine similarity of $l_{2}$-normalized representation pairs.
Following NIST Speaker Recognition evaluation protocol [38], the performance is reported in terms of Equal Error Rate (EER) and minimum Detection Cost Function (minDCF) with $P_{target}=0.01$, $C_{\text{miss}}=1$ and $C_{\text{fa}}=1$.
@@ V-A Effect of the positive sampling on SSL performance
Framework Pos. sampling VoxCeleb1-O $\Delta$ (%)
@@
The supervised baseline achieves an EER of 1.34% on VoxCeleb1-O, highlighting the performance gap that remains between SSL and supervised systems. This work hypothesizes that the primary bottleneck is the same-utterance positive sampling at the core of SSL frameworks. As demonstrated in Table I, significant performance improvements are observed when leveraging VoxCeleb metadata to create anchor-positive pairs from distinct recordings of the same speaker. Supervised positive sampling results in EER reductions of approximately 73%, 72%, 45%, 41%, and 23% for SimCLR, MoCo, SwAV, VICReg, and DINO, respectively, on VoxCeleb1-O.
These findings reveal that the default SSL same-utterance positive sampling strategy is a significant limitation that hinders the performance of SSL frameworks. Adopting a different positive sampling method could significantly enhance the effectiveness of SSL frameworks on SV tasks.
@@ V-B Hyperparameter tuning for SSPS
TABLE II: Hyperparameters search for SSPS-NN and SSPS-Clustering in terms of SV performance. The model is SimCLR, the encoder is Fast ResNet-34, and the benchmark is VoxCeleb1-O. Best results for each method are in bold.
Positive sampling SSPS Hyperparameters VoxCeleb1-O
@@
Figure 4: Hyperparameters search for SSPS-NN (a) and SSPS-Clustering (b) in terms of EER, pseudo-positive speaker accuracy, and pseudo-positive recording accuracy. The model is SimCLR, the encoder is Fast ResNet-34, and the EER is reported on VoxCeleb1-O.
In this work, it is hypothesized that the primary bottleneck of SSL frameworks lies in the way positive samples are selected. The preliminary results of SSPS on SV (VoxCeleb1-O), obtained with SimCLR and the Fast ResNet-34 encoder, with different sampling algorithms and hyperparameters, are reported in Table II. These results are compared with SSL and Supervised positive sampling baselines, achieving 9.41% EER and 3.93% EER on VoxCeleb1-O. Additionally, Figure 4 illustrates the effect of $K$ and $M$ in terms of EER on VoxCeleb1-O, as well as pseudo-positives speaker accuracy and pseudo-positives recording accuracy.